(22 POINTS) 3 ) A block with a mass of 20 kg rests on a hill inclined at 30P to the horizontal. The coefficient of static friction is 0.25 and the coefficient of kinetic friction is 0.15. What is the magnitude of a force (Fa), applied parallel to the incline, needed to move the block UP the hill at a constant velocity? Express your answer in Newtons using 3 significant figures.

help!
Answer #1

プN 20 Kg Fa Fr Mg cos30 Mg

As Shown in FBD of block,-

N = Mg cos30 = (20*10)(0.87) = 173.20N

For moving the block, the friction applied on the box will be kinetic friction force and If we have to keep the velocity constant then acceleration would zero in parallel direction of incline so-

Fnet = Fa - Fr

Ma = Fa - (0.15)N

M*0 = Fa - 0.15(173.20)

0 = Fa - 25.98

Fa = 25.98 N

Thank you !!! Have fun

Please Rate High.